The sides of one triangle measure 3, 4, and 5. The sides of a triangle measure 10, 6, and 8. Are the two triangles similar?

Answer by stanbon(75887)

The sides of one triangle measure 3, 4, and 5. The sides of a triangle measure 10, 6, and 8. Are the two triangles similar?
---
Yes: (3/6)=(4/8)=(5/10)
=================
Cheers,
Stan H.